OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form.

OpenStudy (anonymous):

\[Point-Slope \space Form = y - y_{1} = m(x - x_{1})\]where: \(x_{1}\) = -3 \(y_{1}\) = 4